HIGH-RESOLUTION INFRARED SPECTRA OF BICYCLO[1.1.1]PENTANE: ANALYSES OF THE ν_{14}(e’) AND ν_{18}(a_{2}”) BANDS

M. A. MARTIN and A. PERRY, Department of Chemistry, Oregon State University, Corvallis, OR, 97332-4003, USA; T. MASIELLO, Department of Chemistry and Biochemstry, California State University, East Bay, Hayward, CA, 94542, USA; K. SCHWARTZ, Department of Chemistry, Oregon State University, Corvallis, OR, 97332-4003, USA; J.W. NIBLER, Department of Chemistry, Oregon State University, Corvallis, OR, 97332-4003, USA; \underline{A. WEBER}, Optical Technology Division, National Institute of Standards and Technology,Gaithersburg, MD, 20899-8441, USA; A. MAKI, 15012 24th Ave., S.E., Mill Creek, WA, 98012, USA; T.A. BLAKE, Pacific Northwest National Laboratory, P.O. BOX 999, Mail Stop K8-88, Richland, WA,USA.

Infrared spectra of the ν_{14}(e’) 540 cm^{−1} and ν_{18}(a_{2}”) 835 cm^{−1} bands of bicyclo[1.1.1]pentane have been recorded at a resolution sufficient (0.0015 cm^{−1}) to resolve for the first time individual rovibrational lines. This report presents the ground state constants for this molecule determined from two of the ten infrared-active fundamental bands. The fitted transitions were judged to be free of perturbations of their upper state levels. A combined total of more than 6000 lines with J and K quantum numbers of the two bands ranging up to about 60 were fit with an obs-calc RMS deviation of 0.00021 cm^{−1}. The following principal constants (in units of cm^{−1}) of the rovibrational ground state were obtained: B_{0} = 0.23994136(30), D_{J} = 5.9959(96)x10^{−8}, and D_{JK} = -1.5052(98)x10^{−8}. The band centers were found to be ν_{14} = 540.326397(19) cm^{−1} and ν_{18} = 832.92902(3) cm^{−1}. The values in parentheses represent the uncertainties (2 standard deviations) in the last digits of the value of the constants. The results are compared with those obtained for [1.1.1]propellane and with those computed at the ab initio anharmonic level using the B3LYP density functional method with a cc-pVTZ basis set.